Prove the 3√7 is irrational


1
Expert's answer
2022-05-05T14:43:18-0400

Let us first prove that √7 is an irrational number:

let √7 be rational, then it can be represented as

√7 = p/q - irreducible fraction, where p,q are natural numbers

then

"7=\frac{p^{2}}{q^{2}} , 7q^2=p^2"

Because "7q^2" is divisible by 7, then "p^2" is also divisible by 7,

then p=7k, where k is a natural number.

We get

"7q^2=(7k)^2, 7q^2=49k^2, q^2=7k^2" - so q is divisible by 7.

It turns out that p - is divisible by 7 and q - is divisible by 7, i.e. contradiction,

because p/q is an irreducible fraction. So there is no rational number that is equal to √7.

And accordingly there is no rational number that is equal to 3√7.